How do you solve problems like this?

Probability Mass Function of Geometric (p) distribution is f(x) = (1-p)^x-1 p, x = 1,2,...

If the number of orders this month is a Geom(0,7) random variable, find the probability that we have at most 3 orders.
Answer #1

The following information is provided:-

The population proportion of success is p=0.7, also, 1 - p = 1 - 0.7 = 0.3

We need to compute

Therefore, we get that

This implies that

which means that the probability we are looking for is

Please upvote. Let me know in the comments if anything is unclear. I will reply ASAP!